OpenStudy (anonymous):

What is the equation, in slope-intercept form, of the line that passes through (0, −2) and has a slope of −3? y = −3x − 2 y = −3x + 2 y = −2x + 3 y = −2x − 3 A ?

OpenStudy (anonymous):

@texaschic101 @Olyalhic

OpenStudy (texaschic101):

we will use y = mx + b slope(m) = -3 (0,-2)...x = 0 and y = -2 since we already have the slope(m) and our points (x,y), we just need to find the y intercept, which is b. now we sub y = mx + b -2 = -3(0) + b -2 = b so our equation is : y = -3x - 2

OpenStudy (texaschic101):

do you understand ? Do you have any questions at all ?

OpenStudy (anonymous):

yaaay i was right.. no questions.. ty